OpenStudy (anonymous):

[6.06] Nick works two jobs to pay for college. He tutors for $15 per hour and also works as a bag boy for $8 per hour. Due to his class and study schedule, Nick is only able to work up to 20 hours per week, but must earn at least $150 per week. If t represents the number of hours Nick tutors and b represents the number of hours he works as a bag boy, which system of inequalities represents this scenario? t + b 20 15t + 8b = 150 t + b 20 15t + 8b 150 t + b 20 15t + 8b 150 None of the systems shown represent this scenario.

OpenStudy (agent47):

t+b=20 15t+8b=150

OpenStudy (anonymous):

[6.06] Nick works two jobs to pay for college. He tutors for $15 per hour and also works as a bag boy for $8 per hour. Due to his class and study schedule, Nick is only able to work up to 20 hours per week, but must earn at least $150 per week. If t represents the number of hours Nick tutors and b represents the number of hours he works as a bag boy, which system of inequalities represents this scenario? t + b 20 15t + 8b = 150 t + b 20 15t + 8b 150 t + b 20 15t + 8b 150 None of the systems shown represent this scenario.